Say you are playing Twister and spinning the color wheel. There are four colors: red, green,...

  1. Say you are playing Twister and spinning the color wheel. There are four colors: red, green, blue, and yellow. Spin the spinner 12 times and count blue as a success.

                                                                     

  1. What is the probability of getting the number 3 in one spin (probability of success)?
  2. Assuming this fits all the criteria for a binomial distribution, what would the expected value (mean) be?
  3. What would the standard deviation be?
Answer #1

a) P(3 in one spin) = P(blue) = 1/4 = 0.25

b) n = 12

p = 0.25

Expected value = n * p = 12 * 0.25 = 3

c) Standard deviation = sqrt(n * p * (1 - p)) = sqrt(12 * 0.25 * 0.75) = 1.5